# CONSENT: A Negotiation Framework for Leveraging User Flexibility in Vehicle-to-Building Charging under Uncertainty

### v1

@@ E.3. User Choice Modeling
Figure 7. Hourly distribution of car arrivals and departures over all samples.
@@
Figure 8. Overlay of Initial SoC at Arrival and Required SoC at Departure (bins shifted for clarity).
@@ Appendix F Experiments: More details
To model the piece-wise linear SoC charging curves, we break the charging into 3 segments. The charging power is kept constant at maximum (20 kW) for SoC $\leq$83%, then linearly decreasing as $\frac{4}{3}\cdot\text{SoC}+130$ for 83–90%, and $\text{SoC}+100$ for 90–100%. Discharging assumes constant power with linear SoC decline. Electricity prices follow Silicon Valley Power rates: \$0.178/kWh peak (6 AM–10 PM), \$0.137/kWh off-peak, and an \$11.67/kW demand charge [svp].
F.1. Forecast Generation.
@@
Table 9. Diagnostic metrics for generative EV behavior models
@@
Building Load Forecast. Building demand can usually be predicted to within $\approx$2–10% mean absolute percentage error (MAPE) over day‑ahead horizons[cai2019day, yildiz2017review]. We therefore create a “forecast” by adding zero‑mean Gaussian noise, tuned to that error band, to the realized load. This keeps EV arrivals and departures as the dominant source of uncertainty and avoids entangling them with exogenous load‑forecast errors. A fully coupled forecasting–control stack is left for future work.
@@ Appendix G Results: More Discussion
Table 10. Comparative analysis of different real-time charging policies, with costs scaled for consistency with Table 1. The results highlight the superiority of the MC-MPC strategy, which achieves the lowest average electricity cost among all practical methods, performing within a small margin of the theoretical Oracle (MILP) benchmark. Crucially, MC-MPC accomplishes this economic efficiency without compromising service quality, delivering perfect charging reliability (zero unfulfilled SoC). It also provides the most effective peak shaving, demonstrating a robust balance of cost savings, reliability, and grid stability.
@@
Table 11. Base-case performance of Real‐Time Policies for all months
@@ G.1. MC-MPC as Charging Policy
  • Least Laxity First (LLF): A heuristic that prioritizes charging EVs with the least amount of idle time before their departure [nakahira2017smoothed].
  • Early Deadline First (EDF): A heuristic that prioritizes charging EVs with the earliest departure time [stankovic_EDF].
  • Reinforcement Learning (RL): We use a Deep Deterministic Policy Gradient (DDPG) algorithm to manage charging actions, as detailed in prior work [liu2025]. A separate model is trained for each month.
@@
Table 12. Electricity Cost (USD) of all Policies under Various Uncertainty Scenarios.
@@
Departure‐time uncertainty 7086 $\pm$ 1411 7334 $\pm$ 1480 7357 $\pm$ 1445 7428 $\pm$ 1423 7433 $\pm$ 1420 7534 $\pm$ 1477 8059 $\pm$ 1409
@@
Table 13. Cumulative Missing SoC per month (kWh) over all Policies under Various Uncertainty Scenarios
@@
The RL policy achieves a competitive average cost of \$8808. While it secures lower costs in some summer months by implicitly learning temporal demand patterns, our MC-MPC proves more stable and cost-effective on average. Both intelligent policies significantly outperform the heuristics; MC-MPC is approximately 1.3% more efficient than LLF [nakahira2017smoothed] and EDF [stankovic_EDF]. The gap is even wider when compared to rigid, real-world policies, with our approach outperforming ReqCharge and MaxCharge by 2.2% and 5.4%, respectively.
@@
Table 14. MC-MPC Hyperparameter Search Results
@@
G.2. User Choices and Negotiation Behavior
We aim to understand how simulated users interact with the menu of charging options in our negotiation framework, through Figures 910, and 11.
Figure 9 indicates that most users gravitate toward the exact fulfillment option (Option 0), suggesting a strong preference for maintaining their original charging request without deviation. Nonetheless, while the majority remain inflexible, a substantial subset is willing to accept one of the flexible options (Choices 1, 2, or 3) or even to reject all offers. This highlights the dual reality faced by system designers: exact fulfillment cannot be relied upon exclusively for system flexibility, yet there remains meaningful engagement with negotiated alternatives, providing valuable operational leeway for building operators.
@@ G.2. User Choices and Negotiation Behavior
Figure 10 shows the distribution of departure time deviations across all users. The sharp spike at zero deviation confirms that most users are not shifted from their requested departure, with only small fractions accepting nontrivial delays. Notably, the presence of a secondary group experiencing larger delays underscores that operational flexibility is selectively leveraged: the system offers minimal inconvenience to most, while extracting valuable participation from a subset of more flexible users. The wide tail further emphasizes the heterogeneity in user flexibility and tolerated inconvenience.
Figure 9. Histogram of choice distribution for all users
Figure 10. Histogram of departure deviation for all users
Figure 11. Histogram of SoC deviation for all users
Figure 12. User cost and distribution upon changing the user sensitivity weights using CONSENT
Similarly, Figure 11 shows the distribution of SoC deviations at departure. Note that these deviations are represented as negative values relative to the user’s requested SoC, and thus the data appear primarily in the second quadrant of the plot. Nearly all users experience negligible SoC shortfalls, shown by the pronounced peak near zero deviation. However, isolated secondary peaks at moderate and larger SoC deficits reveal that some negotiated solutions do require a minority of users to accept tangible compromises. The majority’s satisfaction is preserved, while building-level optimization gains cost-effective flexibility by dispersing minor inconveniences among the most adaptable users.
@@
Figure 13. User cost and utility distribution by negotiation option (for users who chose to reject)

### v3

@@ E.3 User Choice Modeling
Figure 8: Hourly distribution of car arrivals and departures over all samples.
@@
Figure 9: Overlay of Initial SoC at Arrival and Required SoC at Departure (bins shifted for clarity).
@@ Appendix F Experiments: More details
To model the piece-wise linear SoC charging curves, we break the charging into 3 segments. The charging power is kept constant at maximum (20 kW) for SoC $\leq$83%, then linearly decreasing as $\frac{4}{3}\cdot\text{SoC}+130$ for 83–90%, and $\text{SoC}+100$ for 90–100%. Discharging assumes constant power with linear SoC decline. Electricity prices follow Silicon Valley Power rates: \$0.178/kWh peak (6 AM–10 PM), \$0.137/kWh off-peak, and an \$11.67/kW demand charge [22].
F.1 Forecast Generation.
@@
Table 9: Diagnostic metrics for generative EV behavior models
@@
Building Load Forecast. Building demand can usually be predicted to within $\approx$2–10% mean absolute percentage error (MAPE) over day‑ahead horizons[6, 30]. We therefore create a “forecast” by adding zero‑mean Gaussian noise, tuned to that error band, to the realized load. This keeps EV arrivals and departures as the dominant source of uncertainty and avoids entangling them with exogenous load‑forecast errors. A fully coupled forecasting–control stack is left for future work.
@@ Appendix G Results: More Discussion
Table 10: Comparative analysis of different real-time charging policies, with costs scaled for consistency with Table 1. The results highlight the superiority of the MC-MPC strategy, which achieves the lowest average electricity cost among all practical methods, performing within a small margin of the theoretical Oracle (MILP) benchmark. Crucially, MC-MPC accomplishes this economic efficiency without compromising service quality, delivering perfect charging reliability (zero unfulfilled SoC). It also provides the most effective peak shaving, demonstrating a robust balance of cost savings, reliability, and grid stability.
@@
Table 11: Base-case performance of Real‐Time Policies for all months
@@ G.1 MC-MPC as Charging Policy
  • Least Laxity First (LLF): A heuristic that prioritizes charging EVs with the least amount of idle time before their departure [17].
  • Early Deadline First (EDF): A heuristic that prioritizes charging EVs with the earliest departure time [23].
  • Reinforcement Learning (RL): We use a Deep Deterministic Policy Gradient (DDPG) algorithm to manage charging actions, as detailed in prior work [13]. A separate model is trained for each month.
@@
Table 12: Electricity Cost (USD) of all Policies under Various Uncertainty Scenarios.
@@
Departure-time uncertainty 7086 $\pm$ 1411 7334 $\pm$ 1480 7357 $\pm$ 1445 7428 $\pm$ 1423 7433 $\pm$ 1420 7534 $\pm$ 1477 8059 $\pm$ 1409
@@
Table 13: Cumulative Missing SoC per month (kWh) over all Policies under Various Uncertainty Scenarios
@@
The RL policy achieves a competitive average cost of \$8808. While it secures lower costs in some summer months by implicitly learning temporal demand patterns, our MC-MPC proves more stable and cost-effective on average. Both intelligent policies significantly outperform the heuristics; MC-MPC is approximately 1.3% more efficient than LLF [17] and EDF [23]. The gap is even wider when compared to rigid, real-world policies, with our approach outperforming ReqCharge and MaxCharge by 2.2% and 5.4%, respectively.
@@
Table 14: MC-MPC Hyperparameter Search Results
@@
G.2 User Choices and Negotiation Behavior
We aim to understand how simulated users interact with the menu of charging options in our negotiation framework, through Figures 10(a)10(b), and 10(c).
Figure 10(a) indicates that most users gravitate toward the exact fulfillment option (Option 0), suggesting a strong preference for maintaining their original charging request without deviation. Nonetheless, while the majority remain inflexible, a substantial subset is willing to accept one of the flexible options (Choices 1, 2, or 3) or even to reject all offers. This highlights the dual reality faced by system designers: exact fulfillment cannot be relied upon exclusively for system flexibility, yet there remains meaningful engagement with negotiated alternatives, providing valuable operational leeway for building operators.
@@ G.2 User Choices and Negotiation Behavior
Figure 10(b) shows the distribution of departure time deviations across all users. The sharp spike at zero deviation confirms that most users are not shifted from their requested departure, with only small fractions accepting nontrivial delays. Notably, the presence of a secondary group experiencing larger delays underscores that operational flexibility is selectively leveraged: the system offers minimal inconvenience to most, while extracting valuable participation from a subset of more flexible users. The wide tail further emphasizes the heterogeneity in user flexibility and tolerated inconvenience.
(a) Choice distribution
(b) Departure deviation
(c) SoC deviation
Figure 10: Histograms of user negotiation behavior: choice distribution, departure time deviation, and SoC deviation for all users.
Figure 11: User cost and distribution upon changing the user sensitivity weights using CONSENT
Similarly, Figure 10(c) shows the distribution of SoC deviations at departure. Note that these deviations are represented as negative values relative to the user’s requested SoC, and thus the data appear primarily in the second quadrant of the plot. Nearly all users experience negligible SoC shortfalls, shown by the pronounced peak near zero deviation. However, isolated secondary peaks at moderate and larger SoC deficits reveal that some negotiated solutions do require a minority of users to accept tangible compromises. The majority’s satisfaction is preserved, while building-level optimization gains cost-effective flexibility by dispersing minor inconveniences among the most adaptable users.
@@
Figure 12: User cost and utility distribution by negotiation option (for users who chose to reject)